OpenStudy (anonymous):

Pleaaase help D: Davids scores on the first three of four 100-point math tests were 86, 92, and 89. What score does he need on his fourth math test to ensure an average score of at least 90? A. x > 89 B. x > 93 C. x > 78 D. x > 90

OpenStudy (anonymous):

D. Find the average of the already given scores. \[\frac{86+92+89}{3} = 89\] David's scores are equivalent to three 89 scores. The wanted score of 90 is equivalent to four 90 scores. x is the score of the test we want. \[3(89)+x=4(90)\] which is\[267+x=360; ~ \therefore~~ x=360-267 \therefore~x=93\] 93 is greater than 90 but not greater than 93.

OpenStudy (anonymous):

(Finding 89 is not necessary; we could have just added 86+92+89)

OpenStudy (anonymous):

(86 + 92 + 89 + x) / 4 >= 90 267 + x >= 360 => x >= 93

OpenStudy (anonymous):

B. x > 93